The Biloxi Company has the following cost structure: fixed costs of $70,000 per month and variable costs of $50 per unit. The Birmingham Company has the following cost structure: fixed costs of $60,000 per month and variable costs of $60 per unit. Both companies make the same product, which sells for $100 per unit. There is a sales level at which these two companies earn the same profits. What is that sales level? Which company is more profitable as sales volume exceeds this sales level?
Answer #1

rate positively ..

Biloxi Company Birmingham Company
Sales price per unit = 100 100
Less Variable cost per unit = 50 60
contribution margin 50 40
Fixed cost 70000 60000
Lets assume unit at break even is x
Therefore x*50-70000=x*40-60000
50x-10x= 10000
x= 1000
Therefore at 1000 unit both company will have the same profit.
If the sales volume exceed the 1000 unit Biloxi Company will make more profit .